# Novelty and Impact of Economics Papers

### v1

@@ Abstract
00footnotetext: We thank Annie Liang, Ben Golub, Jeff Ely, and Jessica Hullman for helpful comments. We also thank the Quest high performance computing facility at Northwestern University for computational resources and staff assistance.
@@ 1 Introduction
A central challenge for researchers, funding agencies, and academic institutions is to identify which new ideas are most likely to generate significant scientific impact. While novelty is widely regarded as a key driver of impact, its complex nature and the mechanisms through which it operates remain poorly understood. This challenge reflects a fundamental tension in research strategy: should efforts focus on work with high uniqueness, characterized by profound originality in questions and methods, or on work with good timing that taps into ”hot” topics currently capturing the scientific community’s attention? Although both uniqueness and timing are recognized as critical dimensions of novelty, existing measures typically fail to disentangle them. Citation-based approaches (Uzzi et al., 2013; Lee et al., 2015; Trapido, 2015) overlook a paper’s semantic content, while keyword-based methods (Azoulay et al., 2011; Mishra and Torvik, 2016; Arts et al., 2025) cannot quantify its departure from prior work.
@@
We first establish a framework for measuring the multidimensional nature of both novelty and impact. Following Kuhn and Hacking (1970), who argued that scientific ideas are embedded in the text of the scientific literature, we propose that novelty can be measured by quantifying a paper’s semantic neighborhood. We decompose this concept into two dimensions. Spatial novelty measures the semantic isolation of a paper from all prior work at its publication date, capturing its intellectual distinctiveness. Temporal novelty measures the rate of semantic change within a paper’s immediate intellectual neighborhood, reflecting its engagement with a dynamic research frontier. The two dimensions describe a paper’s position within the evolving intellectual landscape.
To measure impact, we employ two widely used but conceptually distinct metrics. The first is citation count, which reflects a paper’s popularity and the research community’s collective judgment of its value. The second is the disruption index (Wu et al., 2019), which measures whether a paper’s contribution renders prior work obsolete (disruptive) or serves as a cumulative link between past and future research (consolidating).
@@ 1.1 Related Work
The measurement of scientific novelty has traditionally relied on citation networks and textual data (Zhao and Zhang, 2025). A prominent citation-based approach is the atypicality score, which quantifies novelty based on unconventional combinations of referenced journals (Uzzi et al., 2013). This concept has been extended to account for same-year publications (Lee et al., 2015) and to measure the unusualness of knowledge recombination within a specific field (Trapido, 2015).
Textual data offers a more direct window into a paper’s content. One common approach assumes that new words or word combinations signal novel ideas. For example, novelty has been measured using the average age of MeSH [^1] MeSH (Medical Subject Headings) is the NLM controlled vocabulary thesaurus used for indexing articles in PubMed. terms (Azoulay et al., 2011), their age and reuse frequency (Mishra and Torvik, 2016), and the tracking of new word combinations and their subsequent reuse (Arts et al., 2025). However, keyword-based methods can be biased by polysemy and often overlook deeper semantic relationships. To mitigate this, other work has extracted entities as proxies for knowledge (Luo et al., 2022; Chen et al., 2024) or identified novel contribution sentences using semantic analysis (Wang et al., 2023). Another approach assessed novelty by computing semantic distances between a paper’s references, represented by title embeddings (Shibayama et al., 2021). Our method departs from these proxies by analyzing the full text of a paper and using LLMs to generate comprehensive semantic representations. This approach captures a richer representation of a paper’s content than methods relying solely on titles, abstracts, or keywords. By computing semantic distances between papers, we construct metrics that directly measure a paper’s departure from the existing literature and its engagement in research frontier.
@@
Novelty measures are typically validated through two primary approaches. Direct validation involves correlating proposed metrics with ground-truth labels derived from expert evaluations or peer review outcomes (Tahamtan and Bornmann, 2018; Shibayama et al., 2021). Indirect validation, which is more common, assesses a metric’s correlation with established novelty indices or its predictive power for outcomes like citation counts (Luo et al., 2022; Foster et al., 2021; Shi and Evans, 2023; Arts et al., 2025). We adopt the latter approach, validating our measures by evaluating their ability to predict citation counts and disruption.
@@
Scientific novelty has been conceptualized through various theoretical lenses (Zhao and Zhang, 2025). A foundational dichotomy distinguishes between the introduction of entirely new elements (Berlyne, 1960; Yan et al., 2020; Luo et al., 2022) and the novel recombination of existing elements, a concept rooted in Schumpeterian innovation theory (Schumpeter, 1983; Nelson and Winter, 1985; Guetzkow et al., 2004). Other frameworks conceptualize novelty in terms of network structures, where new connections bridge previously disconnected knowledge communities (Foster et al., 2015; Hofstra et al., 2020), or from a psychological perspective, where novelty is a function of surprise relative to a reader’s cognitive schema (Foster et al., 2021). We use a paper’s position within the intellectual landscape as its novelty. This captures not only the semantic difference between the focal paper and other work, but also the evolvement of the focal paper’s semantic neighborhood.
@@
Prior literature has also sought to create typologies of novelty, generally along two distinct axes. The first stream differentiates novelty by its form or mechanism. This includes distinctions between novelty arising from the age and volume of concepts (Mishra and Torvik, 2016), the recombination of existing knowledge versus the creation of new elements (Yan et al., 2020), or between a paper’s content and its context (Shi and Evans, 2023). A second stream distinguishes novelty by its locus or domain of contribution, such as new theory, new methodology, or new empirical results (Leahey et al., 2023). Our framework contributes to this literature by proposing a new form-based typology—spatial versus temporal novelty—that is conceptually orthogonal to the locus of contribution. A key advantage of our measurement approach is its ability to operationalize this orthogonality. By decomposing each paper into distinct insight types, we can measure the spatial and temporal novelty of a paper’s theory, its methodology, or its core question-conclusion pair, thereby providing a more granular understanding of scientific innovation.
@@
Computational analysis of scientific texts has evolved significantly. Early work relied on sparse vector representations such as term frequency–inverse document frequency (TF–IDF) and probabilistic topic models like Latent Dirichlet Allocation (LDA) (Blei et al., 2003) to map research trends (Gupta et al., 2022; Egger and Yu, 2022). More recent methods leverage dense vector representations (embeddings) to model document semantics more effectively. Embedding-based models like SPECTER (Cohan et al., 2020) and domain-specific word embeddings (Tshitoyan et al., 2019) have improved performance on tasks such as classification, clustering, and citation prediction by capturing nuanced contextual relationships.
Building on these foundations, large language models (LLMs) now enable direct synthesis and representation of scientific text at an unprecedented scale. Current applications include screening for systematic reviews (Dennstädt et al., 2024), generating retrieval-augmented literature surveys (Han et al., 2024), classifying research articles (Raja et al., 2023), and predicting bibliometric outcomes (Vital Jr et al., 2025). LLMs have also proven effective in summarization tasks, producing outputs for both expert and lay audiences (Pakull et al., 2024; Ding et al., 2024) and even drafting review articles (Wang et al., 2024). Our work contributes to this literature by using LLMs not just for summarization but to create the rich semantic representations that underpin our measurement of novelty.
@@ 2.1 Measuring Multidimensional Impact
Our second measure captures a paper’s capacity to redirect a field’s intellectual trajectory. We employ the Disruption Index (DI) developed by Funk and Owen-Smith (2017) and Wu et al. (2019), which we classify as an impact measure rather than a novelty measure Leibel and Bornmann (2024). Whereas citations measure the magnitude of influence, the DI measures its nature: whether a paper’s contribution displaces prior work or builds cumulatively upon it. The index is derived from the citation network surrounding a focal paper (FP) and is defined as:
@@ 2.2 Novelty as Semantic Position
The study of scientific progress hinges on the concept of novelty, yet its measurement remains a central challenge. Prevailing methods, such as the widely-used atypicality score that measures unconventional reference combinations (Uzzi et al., 2013), have provided powerful insights. However, these approaches share a common conceptual foundation: they treat novelty as an intrinsic attribute of a paper, inferred from its inputs (e.g., references) or its internal content (e.g., new keywords). This view, while intuitive, limits our ability to understand the strategic and ecological dynamics of scientific discovery.
@@ 3.1 Data Source and Dataset Construction
Our analysis draws on a comprehensive corpus of scholarly articles in economics. The initial dataset comprises 105,082 articles published between 1980 and 2020, sourced from RePEc, Google Scholar, and SciSciNet (Lin et al., 2023). From SciSciNet, we obtain key bibliometric variables, including the disruption index and the atypicality score. We supplement this with citation counts collected from Google Scholar in December 2024.
@@
Table 1: Descriptive Statistics by Economic Subfield. “Disruptive rate” is the fraction of papers with a disruption index greater than zero, following Wu et al. (2019).
@@ 3.2 Semantic Content Extraction and Representation
Third, to ensure the robustness of our semantic representations, we embed each of the five summary types using three complementary embedding models: SciBERT (Beltagy et al., 2019), a BERT model pre-trained on a large corpus of scientific literature; SPECTER (Cohan et al., 2020), a model specifically designed to generate document-level embeddings for scientific papers; and Qwen (Li et al., 2023), a general-purpose state-of-the-art LLM. This triangulation approach mitigates the risk of model-specific biases.
@@
The use of modern LLMs trained on vast internet corpora introduces a potential for “look-ahead bias,” where the model might inadvertently incorporate information unavailable at the time a paper was written (Sarkar and Vafa, 2024). We mitigate this concern through several strategies. First, our LLM task is summarization, not prediction. Different from prediction task such as predicting the citation count of a paper will receive by 2024 (which is likely to be in a LLM’s training set), in our task the model is prompted to condense the provided text, not to infer future outcomes or context. Second, our prompts are designed to be content-focused, explicitly instructing the model to avoid external contextualization. Third, our metrics are based on relative semantic distances within a historical cohort, making them less sensitive to absolute shifts in the embedding space.
@@ 4.1 Experimental Design
  • Atypicality: The combination-based novelty scores developed by Uzzi et al. (2013), including 10th percentile, median, and number of combinations, as provided by SciSciNet (Lin et al., 2023).
@@ 4.2 Predicting Log Citation Counts
We first assess the ability of our isolation metrics to predict long-term citation counts, a standard proxy for scientific impact. Our primary evaluation metric is the out-of-sample Mean Squared Error (MSE). We also report the ”completeness score” from Fudenberg et al. (2022), which measures the proportional reduction in MSE relative to a baseline feature set (Controls + Subfield). The larger the completeness (maximum is 1), the more error the tested model reduce from the base, which means the better the model’s performance. For example, using all features together gives a completeness score of 34.92%, meaning that this large set of features can reduce 34.92% of our baseline prediction error. This suggests that predicting log citation counts is a hard task.
@@ 4.3 Predicting Disruptive Papers
We next turn to a more nuanced prediction task: identifying disruptive papers. This allows us to test whether our isolation metrics are merely a proxy for future attention (citations) or if they capture a more fundamental aspect of a paper’s contribution. We frame this as a binary classification task, classifying a paper as disruptive if its Disruption Index (DI) is greater than zero (Wu et al., 2019).
@@ 5.3 The Structure of the Novelty Space
While different fields tend to cluster in different quadrants, these mappings reflect only median tendencies and mask considerable within-field heterogeneity. Individual papers within every field span multiple archetypes. For instance, Behavioral Economics — though exhibiting higher novelty on average — includes substantial Consolidating work that refines foundational insights alongside Trailblazing explorations of new domains. This diversity underscores that our framework characterizes research strategies, not research quality: each quadrant represents a legitimate scholarly approach with distinct contributions to knowledge accumulation. Importantly, our framework also reveals substantial overlap across fields with distinct median profiles. Comparing the full distributions of Macroeconomics and Economic History confirms that our novelty scores identify research archetypes that transcend field boundaries. Appendix D provides detailed distributions for all subfields, further substantiating both the structural trade-off and the field-level patterns.
Figure 7: Median novelty profiles by economic subfield. Each point represents the median spatial and temporal novelty score for all papers in a given subfield. Error bars represent the standard error of the mean.
@@ B.1 Dataset Construction
Our dataset is constructed from three primary sources: RePEc (Research Papers in Economics) for metadata, Google Scholar for citation counts and PDF access, and SciSciNet (Lin et al., 2023) for bibliometric indicators. The scope of our corpus is defined by the list of economics journals compiled by Angrist et al. (2017), which includes a broad set of general interest and top field journals.
@@
Table 6: List of Economics Journals included in the Corpus. We use the list of journals from Angrist et al. (2017).
@@ E.1 Robustness of High-Citation Analysis
Novelty Archetype High-Cite Rate Std 95% CI lower 95% CI upper
Consolidating 0.1808 0.0031 0.1748 0.1868
Outlying 0.1970 0.0027 0.1918 0.2023
Trendy 0.2914 0.0028 0.2858 0.2969
Trailblazing 0.3247 0.0032 0.3185 0.3308
all 0.2477 0.0001 0.2475 0.2479
@@
Novelty Archetype High-Cite Rate Std 95% CI lower 95% CI upper
Consolidating 0.0646 0.0021 0.0605 0.0687
Outlying 0.0736 0.0017 0.0702 0.0770
Trendy 0.1178 0.0020 0.1139 0.1216
Trailblazing 0.1345 0.0029 0.1289 0.1401
all 0.0973 0.0001 0.0971 0.0975
@@
Novelty Archetype High-Cite Rate Std 95% CI lower 95% CI upper
Consolidating 0.0054 0.0006 0.0042 0.0065
Outlying 0.0074 0.0005 0.0064 0.0084
Trendy 0.0108 0.0005 0.0098 0.0118
Trailblazing 0.0137 0.0007 0.0124 0.0151
all 0.0093 0.0001 0.0092 0.0094

### v3

@@ Abstract
00footnotetext: We thank Annie Liang, Ben Golub, Jeff Ely, Jessica Hullman, and Anqi Zhang for helpful comments. We also thank the Quest high performance computing facility at Northwestern University for computational resources and staff assistance.
@@ 1 Introduction
A central challenge for researchers, funding agencies, and academic institutions is to identify which new ideas are most likely to generate significant scientific impact. While novelty is widely regarded as a key driver of impact, its complex nature and the mechanisms through which it operates remain poorly understood. This challenge reflects a fundamental tension in research strategy: should efforts focus on work with high uniqueness, characterized by profound originality in questions and methods, or on work with good timing that taps into ”hot” topics currently capturing the scientific community’s attention? Although both uniqueness and timing are recognized as critical dimensions of novelty, existing measures typically fail to disentangle them. Citation-based approaches (uzzi2013atypical; lee2015creativity; trapido2015novelty) overlook a paper’s semantic content, while keyword-based methods (azoulay2011incentives; mishra2016quantifying; arts2025beyond) cannot quantify its departure from prior work.
@@
We first establish a framework for measuring the multidimensional nature of both novelty and impact. Following kuhn1970structure, who argued that scientific ideas are embedded in the text of the scientific literature, we propose that novelty can be measured by quantifying a paper’s semantic neighborhood. We decompose this concept into two dimensions. Spatial novelty measures the semantic isolation of a paper from all prior work at its publication date, capturing its intellectual distinctiveness. Temporal novelty measures the rate of semantic change within a paper’s immediate intellectual neighborhood, reflecting its engagement with a dynamic research frontier. The two dimensions describe a paper’s position within the evolving intellectual landscape.
To measure impact, we employ two widely used but conceptually distinct metrics. The first is citation count, which reflects a paper’s popularity and the research community’s collective judgment of its value. The second is the disruption index (wu2019large), which measures whether a paper’s contribution renders prior work obsolete (disruptive) or serves as a cumulative link between past and future research (consolidating).
@@ 1.1 Related Work
The measurement of scientific novelty has traditionally relied on citation networks and textual data (zhao2025review). A prominent citation-based approach is the atypicality score, which quantifies novelty based on unconventional combinations of referenced journals (uzzi2013atypical). This concept has been extended to account for same-year publications (lee2015creativity) and to measure the unusualness of knowledge recombination within a specific field (trapido2015novelty).
Textual data offers a more direct window into a paper’s content. One common approach assumes that new words or word combinations signal novel ideas. For example, novelty has been measured using the average age of MeSH [^1] MeSH (Medical Subject Headings) is the NLM controlled vocabulary thesaurus used for indexing articles in PubMed. terms (azoulay2011incentives), their age and reuse frequency (mishra2016quantifying), and the tracking of new word combinations and their subsequent reuse (arts2025beyond). However, keyword-based methods can be biased by polysemy and often overlook deeper semantic relationships. To mitigate this, other work has extracted entities as proxies for knowledge (luo2022combination; chen2024exploring) or identified novel contribution sentences using semantic analysis (wang2023measuring). Another approach assessed novelty by computing semantic distances between a paper’s references, represented by title embeddings (shibayama2021measuring). Our method departs from these proxies by analyzing the full text of a paper and using LLMs to generate comprehensive semantic representations. This approach captures a richer representation of a paper’s content than methods relying solely on titles, abstracts, or keywords. By computing semantic distances between papers, we construct metrics that directly measure a paper’s departure from the existing literature and its engagement in research frontier.
@@
Novelty measures are typically validated through two primary approaches. Direct validation involves correlating proposed metrics with ground-truth labels derived from expert evaluations or peer review outcomes (tahamtan2018creativity; shibayama2021measuring). Indirect validation, which is more common, assesses a metric’s correlation with established novelty indices or its predictive power for outcomes like citation counts (luo2022combination; foster2021surprise; shi2023surprising; arts2025beyond). We adopt the latter approach, validating our measures by evaluating their ability to predict citation counts and disruption.
@@
Scientific novelty has been conceptualized through various theoretical lenses (zhao2025review). A foundational dichotomy distinguishes between the introduction of entirely new elements (berlyne1960conflict; yan2020impact; luo2022combination) and the novel recombination of existing elements, a concept rooted in Schumpeterian innovation theory (schumpeter1983theory; nelson1985evolutionary; guetzkow2004originality). Other frameworks conceptualize novelty in terms of network structures, where new connections bridge previously disconnected knowledge communities (foster2015tradition; hofstra2020diversity), or from a psychological perspective, where novelty is a function of surprise relative to a reader’s cognitive schema (foster2021surprise). We use a paper’s position within the intellectual landscape as its novelty. This captures not only the semantic difference between the focal paper and other work, but also the evolvement of the focal paper’s semantic neighborhood.
@@
Prior literature has also sought to create typologies of novelty, generally along two distinct axes. The first stream differentiates novelty by its form or mechanism. This includes distinctions between novelty arising from the age and volume of concepts (mishra2016quantifying), the recombination of existing knowledge versus the creation of new elements (yan2020impact), or between a paper’s content and its context (shi2023surprising). A second stream distinguishes novelty by its locus or domain of contribution, such as new theory, new methodology, or new empirical results (leahey2023types). Our framework contributes to this literature by proposing a new form-based typology—spatial versus temporal novelty—that is conceptually orthogonal to the locus of contribution. A key advantage of our measurement approach is its ability to operationalize this orthogonality. By decomposing each paper into distinct insight types, we can measure the spatial and temporal novelty of a paper’s theory, its methodology, or its core question-conclusion pair, thereby providing a more granular understanding of scientific innovation.
@@
Computational analysis of scientific texts has evolved significantly. Early work relied on sparse vector representations such as term frequency–inverse document frequency (TF–IDF) and probabilistic topic models like Latent Dirichlet Allocation (LDA) (blei2003latent) to map research trends (gupta2022prediction; egger2022topic). More recent methods leverage dense vector representations (embeddings) to model document semantics more effectively. Embedding-based models like SPECTER (cohan2020specter) and domain-specific word embeddings (tshitoyan2019unsupervised) have improved performance on tasks such as classification, clustering, and citation prediction by capturing nuanced contextual relationships.
Building on these foundations, large language models (LLMs) now enable direct synthesis and representation of scientific text at an unprecedented scale. Current applications include screening for systematic reviews (dennstadt2024title), generating retrieval-augmented literature surveys (han2024automating), classifying research articles (raja2023using), and predicting bibliometric outcomes (vital2025predicting). LLMs have also proven effective in summarization tasks, producing outputs for both expert and lay audiences (pakull2024wispermed; ding2024ccsbench) and even drafting review articles (wang2024using). Our work contributes to this literature by using LLMs not just for summarization but to create the rich semantic representations that underpin our measurement of novelty.
@@ 2.1 Measuring Multidimensional Impact
Our second measure captures a paper’s capacity to redirect a field’s intellectual trajectory. We employ the Disruption Index (DI) developed by funk2017dynamic and wu2019large, which we classify as an impact measure rather than a novelty measure Leibel2024disruption. Whereas citations measure the magnitude of influence, the DI measures its nature: whether a paper’s contribution displaces prior work or builds cumulatively upon it. The index is derived from the citation network surrounding a focal paper (FP) and is defined as:
@@ 2.2 Novelty as Semantic Position
The study of scientific progress hinges on the concept of novelty, yet its measurement remains a central challenge. Prevailing methods, such as the widely-used atypicality score that measures unconventional reference combinations (uzzi2013atypical), have provided powerful insights. However, these approaches share a common conceptual foundation: they treat novelty as an intrinsic attribute of a paper, inferred from its inputs (e.g., references) or its internal content (e.g., new keywords). This view, while intuitive, limits our ability to understand the strategic and ecological dynamics of scientific discovery.
@@ 3.1 Data Source and Dataset Construction
Our analysis draws on a comprehensive corpus of scholarly articles in economics. The initial dataset comprises 105,082 articles published between 1980 and 2020, sourced from RePEc, Google Scholar, and SciSciNet (lin2023sciscinet). From SciSciNet, we obtain key bibliometric variables, including the disruption index and the atypicality score. We supplement this with citation counts collected from Google Scholar in December 2024.
@@
Table 1: Descriptive Statistics by Economic Subfield. “Disruptive rate” is the fraction of papers with a disruption index greater than zero, following wu2019large.
@@ 3.2 Semantic Content Extraction and Representation
Third, to ensure the robustness of our semantic representations, we embed each of the five summary types using three complementary embedding models: SciBERT (Beltagy2019SciBERT), a BERT model pre-trained on a large corpus of scientific literature; SPECTER (cohan2020specter), a model specifically designed to generate document-level embeddings for scientific papers; and Qwen (li2023towards), a general-purpose state-of-the-art LLM. This triangulation approach mitigates the risk of model-specific biases.
@@
The use of modern LLMs trained on vast internet corpora introduces a potential for “look-ahead bias,” where the model might inadvertently incorporate information unavailable at the time a paper was written (sarkar2024lookahead). We mitigate this concern through several strategies. First, our LLM task is summarization, not prediction. Different from prediction task such as predicting the citation count of a paper will receive by 2024 (which is likely to be in a LLM’s training set), in our task the model is prompted to condense the provided text, not to infer future outcomes or context. Second, our prompts are designed to be content-focused, explicitly instructing the model to avoid external contextualization. Third, our metrics are based on relative semantic distances within a historical cohort, making them less sensitive to absolute shifts in the embedding space.
@@ 4.1 Experimental Design
  • Atypicality: The combination-based novelty scores developed by uzzi2013atypical, including 10th percentile, median, and number of combinations, as provided by SciSciNet (lin2023sciscinet).
@@ 4.2 Predicting Log Citation Counts
We first assess the ability of our isolation metrics to predict long-term citation counts, a standard proxy for scientific impact. Our primary evaluation metric is the out-of-sample Mean Squared Error (MSE). We also report the ”completeness score” from fudenberg2022measuring, which measures the proportional reduction in MSE relative to a baseline feature set (Controls + Subfield). The larger the completeness (maximum is 1), the more error the tested model reduce from the base, which means the better the model’s performance. For example, using all features together gives a completeness score of 34.92%, meaning that this large set of features can reduce 34.92% of our baseline prediction error. This suggests that predicting log citation counts is a hard task.
@@ 4.3 Predicting Disruptive Papers
We next turn to a more nuanced prediction task: identifying disruptive papers. This allows us to test whether our isolation metrics are merely a proxy for future attention (citations) or if they capture a more fundamental aspect of a paper’s contribution. We frame this as a binary classification task, classifying a paper as disruptive if its Disruption Index (DI) is greater than zero (wu2019large).
@@ 5.3 The Structure of the Novelty Space
While different fields tend to cluster in different quadrants, these mappings reflect only median tendencies and mask considerable within-field heterogeneity. Individual papers within every field span multiple archetypes. For instance, Behavioral Economics — though exhibiting higher novelty on average — includes substantial Consolidating work that refines foundational insights alongside Trailblazing explorations of new domains. This diversity underscores that our framework characterizes research strategies, not research quality: each quadrant represents a legitimate scholarly approach with distinct contributions to knowledge accumulation. Importantly, our framework also reveals substantial overlap across fields with distinct median profiles. Appendix D.2 presents both the inter-quartile range version and the full distributions for Macroeconomics and Economic History, confirming that our novelty scores identify research archetypes that transcend field boundaries, and provides detailed distributions for all subfields that substantiate the structural trade-off and field-level patterns.
Figure 7: Median novelty profiles by economic subfield. Each point represents the median spatial and temporal novelty score for all papers in a given subfield.
@@ B.1 Dataset Construction
Our dataset is constructed from three primary sources: RePEc (Research Papers in Economics) for metadata, Google Scholar for citation counts and PDF access, and SciSciNet (lin2023sciscinet) for bibliometric indicators. The scope of our corpus is defined by the list of economics journals compiled by angrist2017economic, which includes a broad set of general interest and top field journals.
@@
Table 6: List of Economics Journals included in the Corpus. We use the list of journals from angrist2017economic.
@@ D.2 Descriptive Statistics of the Novelty Framework
Figure 16: Median novelty profiles by economic subfield. Each point represents the median spatial and temporal novelty score for all papers in a given subfield. Error bars represent 25-75 inter-quartile range.
@@ E.1 Robustness of High-Citation Analysis
Neighborhood Archetype High-Cite Rate SE 95% CI lower 95% CI upper
Consolidating 0.1808 0.00031 0.18023 0.18143
Outlying 0.1970 0.00027 0.19650 0.19755
Trendy 0.2914 0.00028 0.29080 0.29191
Trailblazing 0.3247 0.00032 0.32403 0.32527
all 0.2477 0.00001 0.24771 0.24775
@@
Neighborhood Archetype High-Cite Rate SE 95% CI lower 95% CI upper
Consolidating 0.0646 0.00021 0.06416 0.06499
Outlying 0.0736 0.00017 0.07322 0.07390
Trendy 0.1178 0.00020 0.11739 0.11815
Trailblazing 0.1345 0.00029 0.13394 0.13506
all 0.0973 0.00001 0.09725 0.09729
@@
Neighborhood Archetype High-Cite Rate SE 95% CI lower 95% CI upper
Consolidating 0.0054 0.00006 0.00526 0.00549
Outlying 0.0074 0.00005 0.00733 0.00753
Trendy 0.0108 0.00005 0.01071 0.01091
Trailblazing 0.0137 0.00007 0.01359 0.01387
all 0.0093 0.00001 0.00929 0.00931